1. The table given below summarizes the 2019 income statement and end-year balance sheet of Drake’s Bowling Alleys. Drake’s financial manager forecasts a 10% increase in sales and costs in 2020. The ratio of sales to average assets is expected to remain at 0.40. Interest is forecasted at 5% of debt at the start of the year. At the end of 2018 debt was $2,400,000 and assets were $6,960,000. (10 points) Income Statement $ in thousands Sales $ 2,900 (40% of average assets) Costs 2,175 (75% of sales) Interest 120 (5% of debt at start of year) Pretax profit 605 Tax 242 (40% of pretax profit) Net income $ 363 Balance Sheet $ in thousands Net assets $ 7,540 Debt $ 2,400 Equity 5,140 Total $ 7,540 Total $ 7,540 a. What is the expected level of assets at the end of 2020? b. If the company pays out 50% of net income as dividends, how much cash will Drake need to raise in the capital markets in 2020? Assumes debt remains constant. c. If Drake is unwilling to make an equity issue, what will be the debt ratio at the end of 2020?
Answer #1

a. Level of sales in 2020 = 2900 *110% = 3190

Let x be level of assets in 2020

3190 = 40% * ( 7540 + x) /2

x =8410

Expected level of assets = 8410

b.

Sales                         3,190
Costs                         2,393 [75 % of sales]
Interest                             120 [2400*5%]
Pre tax profit                             678
Tax                             271 [40% of pretax profit]
Net income                             407
Less: Dividends paid                             203 [50% of net income]
Income to stockholders                             203
Assets 8410 Debt 2400
Existing equity 5140
Net income 203
New equity 667
8410 8410

Amount to be raised in capital markets = 667

c) If it is unwilling to make then it has to raise through debt. Then debt will be = 2400 + 667 = 3067

Debt ratio = 3067/ 8410 = 36.47%